# Two Views Are Better than One: Monocular 3D Pose Estimation with Multiview Consistency

### v2

@@ Abstract
Deducing a 3D human pose from a single 2D image or 2D keypoints is inherently challenging, given the fundamental ambiguity wherein multiple 3D poses can correspond to the same 2D representation. The acquisition of 3D data, while invaluable for resolving the pose ambiguity, is expensive and requires an intricate setup, often restricting its applicability to controlled lab environments. We improve the performance of deep learning-based monocular human pose estimation models by using multiview data but only during training. To do so we propose a novel loss function, consistency loss, that operates on two views and enables adding additional training data with only 2D supervision without knowing the camera’s intrinsic or extrinsic parameters. This loss enforces that the inferred 3D poses of an entire sequence from one view align with the inferred 3D poses from another view under a similarity transform. Our consistency loss substantially improves performance for fine-tuning with no available 3D data. Furthermore, we demonstrate that using our consistency loss can yield state-of-the-art performance when training models from scratch in a semi-supervised manner. Our experiments demonstrate that just two views offset by 90 degrees are enough to obtain good performance, with only marginal improvements by adding more views. Thus, we enable easy acquisition of domain-specific data by capturing activities with off-the-shelf cameras, eliminating the need for elaborate calibration procedures. This research opens new possibilities for domain adaptation in 3D pose estimation, providing a practical and cost-effective solution to customize models for specific applications. The used dataset, featuring additional views, will be made publicly available.
Keywords:
3D human pose consistency loss monocular 3D pose semi-supervised
Figure 1: We improve monocular performance by applying our consistency loss during training to predicted 3D pose sequences from two different views. The consistency loss penalizes variations between the two predicted pose sequences of the same activity. Note that we only use multiple views during training. For every predicted 3D pose sequence obtained from View A and View B, we compute a similarity transform with Procrustes Analysis. This transformation aligns the predicted poses in sequence A with sequence B. The consistency loss is the average distance in 3D between the two pose sequences post-alignment, shown as dashed red lines. Using Procrustes analysis for this transformation enables us to use cameras with unknown intrinsics and extrinsics.
@@ 1 Introduction
Inferring a 3D human pose from a single 2D image or 2D keypoints is fundamentally an ambiguous problem, with a multiple of 3D poses having identical 2D representations. It is, however, possible to learn a mapping from a single 2D image to a 3D pose. Many methods rely either on 3D training data or on two or more views with a calibrated camera setup for training. We show how 3D pose can be accurately learned using only two synchronized uncalibrated cameras during training and still just a single 2D image for inference.
@@
With this inherent ambiguity, models inferring 3D poses need an underlying representation of how the body can move and, ideally, which movements it can expect to see. To introduce such a prior in the model, it has become popular to rely on large foundation models [46], which have learned how the body moves in general. Even so, these do not always generalize to new movements or applications. This introduces the need for fine-tuning the foundation models to accommodate the less frequently seen movements specific for that domain [46].
@@
Methods for adapting 3D human pose models to different domains and movements have traditionally relied on the availability of new 3D data [44, 18, 45]. However, it can be costly and may not be feasible to set up systems for capturing 3D data in the new domain. Alternative methods have been explored to address these challenges. These methods have demonstrated that 3D pose models can be fine-tuned using 2D data, as suggested by previous work [2]. This fine-tuning process ensures that the inferred 3D joint positions align with 2D keypoints in an image, which can be obtained accurately using readily available methods [16, 42]. Unfortunately, relying solely on 2D supervision from a single view can cause models to forget their depth representation, which is problematic since we want to obtain an accurate 3D representation [11].
To advance fine-tuning with purely 2D supervision, we introduce a new loss function that enforces multiview consistency by requiring the inferred 3D pose from one view to be similar to the inferred 3D pose from another view under a similarity transform. Fig. 1 sketches the concept of our consistency loss.
@@
While we focus on increasing the performance of monocular models and not models utilizing multiple views, many datasets used for training monocular models have multiple views of the scene available [12, 31, 29, 17]. We exploit this by utilizing the multiple views in the SportsPose [12] and Human3.6M [13] datasets to evaluate the feasibility of our loss. We demonstrate the improvements that can be obtained by using multiple views during training while only using a single view at inference time. Additionally, we investigate how many views are necessary to obtain improvements in 3D predictions. This study uses the SportsPose dataset  [12], to demonstrate the performance of our proposed loss for fine-tuning a model to a model. The authors provided us access to all seven views of this dataset, which will be released together with this paper. The SportsPose dataset features complex and challenging sports scenarios, making it an ideal test bed for our domain-adaptive approach. The new views are available on our website [^1] Link removed for anonymity.. Additionally, we demonstrate that our consistency loss can enhance model performance when training from scratch in a semi-supervised manner on the Human3.6M dataset. By utilizing our consistency loss, we obtain state-of-the-art among semi-supervised methods.
Moreover, it resolves the inherent ambiguity associated with fine-tuning on 2D data. When using 2D data for training, the model can be confounded by multiple 3D points that project to the same 2D coordinates. Our multiview consistency loss provides a robust solution to this challenge, substantially enhancing the accuracy and reliability of monocular 3D pose estimation.
@@
Our contributions extend beyond introducing the view-consistency loss for domain-adaptive 3D pose estimation. We also present the first set of baseline results on the SportsPose dataset, demonstrating the effectiveness of our approach. We illustrate how our method enhances 3D pose estimation accuracy in dynamic and complex environments by showcasing a model fine-tuned on the SportsPose dataset. This research opens up new possibilities for domain adaptation in 3D pose estimation, providing a practical and cost-effective solution to customize models for specific applications, and unlocks the possibility of increasing the state-of-the-art in monocular pose estimation by training on large amounts of two-view in the wild data.
@@ 2.1 Monocular 3D human pose models
Two primary approaches exist in the monocular 3D human pose estimation domain. One focuses solely on determining the 3D locations of the human skeleton [33, 44, 38], while the other includes estimating the body shape [43, 2, 21, 18, 24]. The latter category often employs parametric body models such as SMPL [27], which describes the body through shape parameters and pose parameters. Notably, our proposed loss remains applicable to both approaches, as 3D joint positions, as well as shape and pose parameters, can be penalized based on the variation between multiple views.
Irrespective of whether the goal is to estimate pose alone or to include shape parameters, monocular 3D human pose estimation commonly adopts either a one-stage or a two-stage approach. In one-stage approaches, the estimation is directly derived from an image or video input, while two-stage approaches involve lifting estimated 2D poses to 3D space. State-of-the-art monocular models that employ the two-stage approach, lifting 2D poses to 3D, achieve remarkable mean-per-joint precision errors (MPJPE [11]). They reach as low as 17mm [46] when lifting ground truth 2D poses on the Human3.6M dataset [13], and 37mm when lifting estimated 2D poses [46].
@@ 2.2 Multiview 3D human pose models and datasets
Multiple synchronized and calibrated cameras have been extensively used in human pose estimation work [15, 4, 34]. Utilizing calibrated camera setups in such approaches has yielded impressive results, even generating state-of-the-art 3D human pose datasets [12, 31, 29, 17]. These datasets, in turn, play a pivotal role in training and advancing monocular models. However, the practical implementation of multi-camera setups involves intricate calibration and synchronization processes, which often confines data to be collected in controlled laboratory environments.
Approaches that require limited or no 3D supervision have also been explored. Some of these are unsupervised and train on single images by lifting a 2D pose to 3D followed by a random rotation and re-projection to 2D [6, 3, 41, 10]. Others more similar to ours use multiple views of the same person [14, 8, 30, 26, 20]. Liu et al. [26] require a fully calibrated camera setup to predict poses. Others do not require known extrinsics but estimate relative camera poses by decomposing the essential matrix estimated from 2D poses predicted in multiple views. Then, the 3D poses are triangulated using the estimated relative poses, which are then used as training data [20, 8]. Mitra et al. [30] add additional training data from multiview images and use metric learning to enforce that images of the same pose have similar embeddings. The approach of Iqbal et al. [14] is most similar to ours, but they require known camera intrinsic, and can only predict scale normalized poses. They infer 3D poses with a monocular model from multiple views and align them rigidly. During training, they penalize the model for differences in the predicted poses. However, the latter two approaches apply only multiview consistency on single image pairs and not sequences, thus significantly limiting their potential use.
@@ 3 Multiview consistency loss
Instead of relying on a known intrinsic or extrinsic calibration, our consistency loss can be deployed without any prior information about the cameras. Instead, the consistency loss applies a similarity transformation and penalizes differences in the poses of two sequences of the same activity, see Figure 1. Avoiding camera calibration simplifies the training pipeline and gives an efficient alternative for handling data from multiple views.
@@
To compute the scaling, rotation, and translation used to transform $\hat{J}_{a,i}$, we estimate the optimal parameters $\hat{\theta}_{ab}$, as in Equation 3. Here, it should be noted that contrary to how similarity transformations traditionally are computed in 3D human pose estimation to calculate the Procrustes aligned MPJPE, we only compute one transformation, $\hat{\theta}_{ab}$, for the entire sequence and not one per frame as in the PA-MPJPE metric [11].
@@ 4 Experiments
4.1 Test protocol
To evaluate our proposed consistency loss, as defined in Equation 1, we conduct experiments using the SportsPose dataset [12]. Since the original paper does not provide a specified test protocol, we employ a test protocol inspired by Human3.6M [13], wherein subjects are distributed across sets to ensure that no subject appears in the same set.
@@ 4.2 Implementation
While our view-consistency loss is versatile and applicable to any monocular 3D human pose method, we choose to adapt the MotionBERT model by Zhu et al. [46] and fine-tuning it to the SportsPose [12] dataset.
To use SportsPose [12] with MotionBERT [46] we need to preprocess the data first. We convert the definition of body keypoints from COCO [25] keypoints used in SportsPose to the Human3.6M [13] body keypoint format. The keypoints are further transformed from meters to millimeters and, by using the extrinsic camera parameters, transformed from the world coordinate system to each camera. Then, following the approach in [5], the camera coordinates are transformed to pixel coordinates and scaled to be within the range $[-1;1]$.
For the fine-tuning of MotionBERT [46], we employ the weights provided for the DSTformer with a depth of five and eight heads. The sequence length is 243, and the feature and embedding sizes are 512. Adhering to the training protocol suggested by Zhu et al. [46], we fine-tune the models for 30 epochs, using a learning rate of 0.0002 and utilizing the Adam optimizer [19].
@@ 4.3 Fine-tuning with 3D data
To compare to the situation when 3D data is available, we also experiment with fine-tuning with 3D data. We implement the proposed fine-tuning configuration from MotionBERT. This involves using a positional loss, $\mathcal{L}_{\text{pos}}$ directly on the 3D poses, coupled with losses on joint velocities, $\mathcal{L}_{\text{vel}}$, and scale only loss, $\mathcal{L}_{\text{scale}}$, as suggested by Rhodin et al. [35]. This combination results in the combined loss for 3D data,
@@
Table 1: Results on SportsPose [12]. Baseline is MotionBERT [46], which is then fine-tuned with either 2D ($\mathcal{L}_{\text{2D}}$) or 3D ($\mathcal{L}_{\text{3D}}$) supervision with and without our proposed multiview consistency loss $\mathcal{L}_{\text{con}}$. All results in mm, lower is better. MPJPE is mean per joint precision error, and PA is Procrustes aligned MPJPE. All results use predicted 2D poses [16]. Bold is the best performance with only 2D data and bold gray is best performance with 3D ground truth. The two views are in Figure 2. The consistency loss improves performance in both cases but substantially more when 3D supervision is not used.
@@
$\mathcal{L}_{\text{2D}_{\text{con}}}$8, Ours 36.6 22.5 34.2 22.2 41.7 25.2 37.3 23.7 32.0 22.7 36.4 22.0
@@
4.4 Fine-tuning without available 3D data
Refining a model through fine-tuning is still possible when ground truth 3D joint data are unavailable. This fine-tuning involves reprojecting the predicted 3D points onto the image and assessing how well these reprojected keypoints aligns with the ground truth 2D keypoints. However, obtaining precise ground truth 2D poses, although less challenging than gathering 3D poses and not requiring specialized hardware, requires substantial effort and manual annotation.
In practice, the use of ground truth 2D poses is limited due to annotation challenges. Instead, models frequently leverage estimated 2D keypoints from detectors such as HRNet [40], AlphaPose [7] or RTMPose [16] for supervision. Notably, when employing estimated keypoints, the reprojection error is often weighted by the confidence scores provided by the 2D keypoint detectors. This strategy minimizes the impact of less reliable keypoints on the overall training process while maintaining the essential guidance for model refinement.
To validate the performance of our proposed consistency loss, we use RTMPose [16] to predict 2D poses for all frames in the SportsPose dataset [12]. The predicted 2D poses are preprocessed as described in Section 4 and then used to fine-tune the MotionBERT [46] model. When fine-tuning the model without the consistency loss, we solely use the 2D reprojection loss,

### v3

@@ Abstract
Deducing a 3D human pose from a single 2D image is inherently challenging because multiple 3D poses can correspond to the same 2D representation. 3D data can resolve this pose ambiguity, but it is expensive to record and requires an intricate setup that is often restricted to controlled lab environments. We propose a method that improves the performance of deep learning-based monocular 3D human pose estimation models by using multiview data only during training, but not during inference. We introduce a novel loss function, consistency loss, which operates on two synchronized views. This approach is simpler than previous models that require 3D ground truth or intrinsic and extrinsic camera parameters. Our consistency loss penalizes differences in two pose sequences after rigid alignment. We also demonstrate that our consistency loss substantially improves performance for fine-tuning without requiring 3D data. Furthermore, we show that using our consistency loss can yield state-of-the-art performance when training models from scratch in a semi-supervised manner. Our findings provide a simple way to capture new data, e.g. in a new domain. This data can be added using off-the-shelf cameras with no calibration requirements. We make all our code and data publicly available.
Figure 1: We improve monocular performance by applying our consistency loss during training to predicted 3D pose sequences from two different views. The consistency loss penalizes variations between the two predicted pose sequences of the same activity. Note that we only use multiple views during training. For every predicted 3D pose sequence obtained from View A and View B, we compute a similarity transform with Procrustes Analysis. This transformation aligns the predicted poses in Sequence A with Sequence B. The consistency loss is the average 3D distance between the two pose sequences post-alignment, shown as dashed red lines. Using Procrustes analysis for this transformation enables us to use cameras with unknown intrinsics and extrinsics.
@@ 1 Introduction
Inferring a 3D human pose from a single 2D image or 2D keypoints is an ill-posed problem, where the 3D pose is not uniquely defined. The consequence is that most existing models inaccurately predict the depth of the keypoints [11]. It is, however, possible to learn a consistent mapping from a single 2D image to a 3D pose [14, 55]. Previous methods train on 3D data [55] or on two or more views with a calibrated camera setup [14]. Such data is expensive and restricts the settings where data can be recorded. Instead, we use a much simpler setup with two synchronized stationary or moving cameras. Here, the 3D pose in each frame is identical up to a transformation, and in the stationary case, this transformation is the same for all the frames recorded by the two cameras. We estimate this transformation using the Procrustes algorithm and optimize for consistent 3D pose predictions during training. At inference, we use a single 2D image, and our approach leads to state-of-the-art performance. Fig. 1 sketches the concept of our consistency loss.
@@
With this inherent ambiguity, models inferring 3D poses need a representation of how the body can move. It has become popular to rely on large foundation models [55], which have learned how the body moves in general based on supervised learning. Even so, foundation models do not always generalize to new movements in specific applications such as sports. This introduces the need for fine-tuning the foundation models to accommodate the less frequently seen movements specific to that domain [55], for which our model is particularly useful since collecting the data is inexpensive.
@@
Methods for adapting 3D human pose models to different domains and movements have traditionally relied on the availability of new 3D data [52, 18, 53]. However, it can be costly and may not be feasible to set up systems for capturing 3D data in the new domain. Alternative methods have been explored to address these challenges. These methods have demonstrated that 3D pose models can be fine-tuned using 2D data, as suggested by previous work [2]. This fine-tuning process ensures that the inferred 3D joint positions align with 2D keypoints in an image, which can be obtained accurately using readily available methods [16, 49]. Unfortunately, state-of-the-art models based solely on 2D supervision from a single view have too inaccurate depth predictions to be used in actual applications such as sports [11]. Thus, there is a need to advance 3D pose prediction from 2D images.
@@
While we focus on increasing the performance of monocular models and not models utilizing multiple views, many datasets used for training monocular models have multiple views of the scene available [12, 33, 31, 17]. We use the multiple views in the SportsPose [12], Human3.6M [13], and SkiPose [44] datasets to evaluate our approach. We use multiple views during training while only using a single view at inference. Additionally, we investigate how many views are necessary to obtain improvements in 3D predictions. Specifically, we use the SportsPose dataset [12] to investigate the performance using varying numbers of views. We have gained access to all seven views of this dataset, and the original authors have allowed us to release the additional data as part of this paper. The SportsPose dataset features complex and challenging sports scenarios, enabling us to test poses that are rare in other datasets. The new views are available on our website [^1] http://christianingwersen.github.io/SportsPose. Additionally, we test our consistency loss on the Human3.6M dataset trained semi-supervised, resulting in superior performance. By utilizing our consistency loss, we obtain state-of-the-art among semi-supervised methods.
@@
Our contributions extend beyond introducing the view-consistency loss for domain-adaptive 3D pose estimation. We also present the first set of baseline results on the SportsPose dataset, demonstrating the effectiveness of our approach. We illustrate how our method enhances 3D pose estimation accuracy in dynamic and complex environments by showcasing a model fine-tuned on the SportsPose dataset. This research opens new possibilities for domain adaptation in 3D pose estimation, providing a practical and cost-effective solution to customize models for specific applications, and unlocks the possibility of increasing the state-of-the-art accuracy in monocular pose estimation by training on large amounts of two-view in-the-wild data.
@@ 2.1 Monocular 3D human pose models
There are two primary approaches for monocular 3D human pose estimation. One solely predicts the 3D locations of the human skeleton [36, 52, 42], while the other includes estimating the body shape [50, 2, 24, 18, 28]. Most models that include the body shape employ parametric body models such as SMPL [30], which describes the body through shape parameters and pose parameters. Notably, our proposed loss remains applicable to both approaches, as 3D joint positions, as well as shape and pose parameters can be penalized based on the variation between multiple views.
Irrespective of whether the goal is to estimate pose alone or to include shape parameters, monocular 3D human pose estimation commonly adopts either a one-stage or a two-stage approach. In the one-stage approaches, the estimation is directly derived from an image or video input, while the two-stage approaches involve lifting estimated 2D poses to 3D space. State-of-the-art monocular models that employ the two-stage approach, lifting 2D poses to 3D, achieve mean-per-joint precision errors (MPJPE) as low as 17mm [55] when lifting ground truth 2D poses on the Human3.6M dataset [13], and 37mm when lifting estimated 2D poses [55]. However, when the same methods are evaluated on other datasets they have much higher MPJPE, it is clear that further work is still required [11].
@@ 2.2 Multiview 3D human pose models and datasets
Multiple synchronized and calibrated cameras have been extensively used to generate data to develop human pose estimation models [15, 4, 38]. Utilizing calibrated camera setups in such approaches has yielded impressive results, and has also been the basis for generating state-of-the-art 3D human pose datasets [12, 33, 31, 17]. These datasets have been essential for developing monocular pose models. However, the practical implementation of multi-camera setups involves a calibrated camera setup. Thus, most data has been collected in controlled laboratory environments, which does not reflect data variability in many scenarios.
Approaches that require limited or no 3D supervision have also been explored. Some of these are unsupervised and train on single images by lifting a 2D pose to 3D followed by a random rotation and re-projection to 2D [5, 3, 48, 10]. Others use multiple views of the same person [14, 6, 32, 29, 23], which is close to our approach. However, these approaches require camera calibration [29]. Some methods only rely on intrinsic calibration and estimate relative camera poses by decomposing the essential matrix estimated from 2D poses predicted in multiple views. Then, the 3D poses are triangulated using the estimated relative poses and used as training data [23, 6]. Mitra et al. [32] add additional training data from multiview images and use metric learning to enforce that images of the same pose have similar embeddings. Similar to our approach, Iqbal et al. [14] rigidly align 3D poses predicted from multiple views but opposed to our method they require known camera intrinsic and can only predict scale normalized poses. During training, they penalize the model for differences in the predicted poses. Both Iqbal et al. [14] and Mitra et al. [32] apply multiview consistency on single image pairs and not sequences, thus risk reducing accuracy and potential use by ignoring temporal information.
@@ 3 Multiview consistency loss
In our approach, we propose a loss function that optimizes for consistency between a sequence of 3D poses predicted from multiple views. The consistency is measured as the distance between Procrustes aligned pose sequences. Therefore, we do not need to know the cameras’ intrinsic or extrinsic calibration or other prior information. Instead, the consistency loss applies a similarity transformation and penalizes differences in the poses of two sequences of the same activity, see Figure 1. Avoiding camera calibration simplifies the training pipeline and gives an efficient alternative for handling data from multiple views.
@@
To compute the scaling, rotation, and translation used to transform $\hat{J}_{a,i}$, we estimate the optimal parameters $\hat{\theta}_{ab}$, as in Equation 3. Here, it should be noted that contrary to how similarity transformations are traditionally computed in 3D human pose estimation to calculate the Procrustes aligned MPJPE, we only compute one transformation, $\hat{\theta}_{ab}$, for the entire sequence and not one per frame as in the PA-MPJPE metric [11]
@@ 4 Experiments
We conduct experiments with fine-tuning a pretrained pose estimator on a new dataset with ground truth 3D (Sec. 4.2) and without any 3D data (Sec. 4.3). In Sec. 4.4 we show the applicability of our loss when training from scratch in a weakly supervised setting. Finally, we investigate the number of views our consistency loss needs (Secs. 4.5 and 4.6).
Datasets
To evaluate our method we have chosen several datasets that all contain multiple views and ground truth 3D poses.
Human3.6M [13] is the most widely used dataset in human pose estimation. It is recorded with four fixed cameras and a total of 3.6 million frames across seven subjects.
SkiPose [44, 40] is a domain-specific dataset of 12 sequences of alpine skiers recorded with six PTZ cameras with a total of 10,197 frames. This lets us test the performance of our loss in the challenging setting of non-fixed cameras.
SportsPose contains several movements common in sports that pretrained models struggle with [12]. It is recorded with seven fixed cameras and has a total of 1.5 million frames, which lets us test the impact of the number of views on our loss.
SportsPose test protocol
As Ingwersen et al. [12] do not provide a specified test protocol, we employ a test protocol inspired by Human3.6M [13], wherein subjects are distributed across sets to ensure that no subject appears in the same set.
@@ 4.1 Implementation
While our consistency loss is versatile and applicable to any monocular 3D human pose method, we choose to adapt the MotionBERT model by Zhu et al. [55] due to its impressive performance on multiple datasets. For fine-tuning the SportsPose [12] dataset we used 2D poses predicted by RTMPose [16].
Details of the preprocessing of the detected 2D poses are in the supplementary material.
For the fine-tuning of MotionBERT [55], we employ the weights provided for the DSTformer with a depth of five and eight heads. The sequence length is 243, and the feature and embedding sizes are 512 as in the original paper. Adhering to the training protocol suggested by Zhu et al. [55], we fine-tune the models for 30 epochs, using a learning rate of 0.0002 and utilizing the Adam optimizer [22].
@@ 4.2 Fine-tuning with 3D data
To compare to the situation when 3D data is available, we also experiment with fine-tuning with 3D data. We implement the proposed fine-tuning configuration from MotionBERT. This involves using a positional loss, $\mathcal{L}_{\text{pos}}$, directly on the 3D poses, coupled with losses on joint velocities, $\mathcal{L}_{\text{vel}}$, and scale only loss, $\mathcal{L}_{\text{scale}}$, as suggested by Rhodin et al. [39]. This combination results in the combined loss for 3D data,
@@
Table 1: Results on SportsPose [12]. Baseline is MotionBERT [55], which is then fine-tuned with either 2D ($\mathcal{L}_{\text{2D}}$) or 3D ($\mathcal{L}_{\text{3D}}$) supervision with and without our proposed multiview consistency loss $\mathcal{L}_{\text{con}}$. All results are in mm where lower is better. MPJPE is mean per joint precision error, and PA is Procrustes aligned MPJPE. All results use predicted 2D poses [16]. Bold is the best performance with only 2D data and bold gray is best performance with 3D ground truth. The two views are shown in Figure 2. The consistency loss improves performance for both 2D and 3D but substantially more when 3D supervision is not used.
@@
$\mathcal{L}_{\text{2D}_{\text{con}}^{\text{frame}}}$, Ours 33.9 21.9 31.0 20.1 36.1 23.0 36.5 23.9 34.1 23.5 34.4 22.5
$\mathcal{L}_{\text{2D}_{\text{con}}}$8, Ours 35.4 20.9 36.2 22.7 40.9 26.1 33.5 22.6 35.1 21.5 36.2 22.8
Table 2: Results on SkiPose [44] Methods in gray use all six views during inference, the rest are monocular. "F" indicates fine-tuning a pretrained model, "S" is self supervised using only 2D on SkiPose. Best performance without ground truth 2D annotations is marked in bold.
Method MPJPE PMPJPE Uses
Metapose[45] - $42$ S
Zhou et al. [54] $42.2$ $29.4$ F+3D
MotionBERT [55] $259$ $132$
CanonPose: [47] $128.1$ $89.6$ S
MHCanonNet [21] $122$ $50.7$ S
Yang et al. [51] - $68.4$ S
Kim et al. [20] $115.2$ $78.8$ S
Pavllo et al. [35]3 $106$ $88.1$ F+S
PoseAug [8] [^3] Value for [8] and [35] reported by [7]. $105.4$ $83.5$ F+S
AdaptPose[7] $99.4$ $83.0$ F+S
$\mathcal{L}_{\text{2D}_{\text{con}}^{\text{frame}}}$, Ours $62.0$ $40.5$ F+S
[…]
@@
4.3 Fine-tuning without 3D data
Our consistency loss is especially useful when fine-tuning a model on a new dataset where multiple views are available but 3D poses are not.
The common way of utilizing additional data without 3D poses during training is to penalize deviations between the 2D poses and the reprojection of the predicted 3D poses [19]. When this is done with ground truth 2D it can work, but it can result in poor results due to the 2D-3D ambiguity [11]. Obtaining ground-truth 2D poses further requires manual annotations. Instead, predicted 2D poses are commonly used. We denote this 2D reprojection loss as